If the probability of A union B is 2/3, the probability of A intersection B is 1/3, and the probability of A is 1/2, then

a.

the probability of B is 5/6

b.

the probability of A given B is 3/5

c.

A and B are not independent

d.

All of the above
Answer #1

P(A union B) = P(A) +P(B) - P(A intersection B)

2/3 = 1/2 + P(B) - 1/3

2/3 -1/2 +1/3 = P(B)

P(B) =( 4-3 +2)/6 = 3/6 = 1/2.

P(A intersection B) = P(B) P(A/B)

1/3 = 1/2 P(A/B)

P(A/B) = 2/3

Probability of A given B = 2/3

A and B are not independent because if A and B are independent ,then P(A intersection B) should be equal to P(A) multiplied by P(B).

Hence, option(C) is correct.